### v1

@@ 2.3 Mixture of Experts
Mixture of Experts (MoE) is a classical ensemble learning framework that combines multiple models to achieve more expressive and accurate predictions (Jacobs et al., 1991; Jordan and Jacobs, 1994). Formally, an MoE model consists of $N$ expert networks, denoted by $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$ for $i=1,\dots,N$. A gating function, $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$, then dynamically determines the contribution of each expert for a given input ${\bm{x}}\in\mathbb{R}^{d}$. The gating function depends on a set of learned score functions, $s_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, associated with each expert, resulting in the following formulation:
${\mathbf{y}}$ $=\sum_{j=1}^{N}G({\bm{x}})_{j}\cdot f_{j}({\bm{x}})=\sum_{j=1}^{N}\frac{\exp\left(s_{j}({\bm{x}})\right)}{\sum_{\ell=1}^{N}\exp\left(s_{\ell}({\bm{x}})\right)}\cdot f_{j}({\bm{x}}),$
where $G(\bm{x})=\mathrm{softmax}(s_{1}(\bm{x}),...,s_{N}(\bm{x}))$. Building on this concept, Shazeer et al. (2017) proposed Sparse Mixture of Experts (SMoE) architecture to efficiently scale up large models. This is achieved by utilizing a sparse gating function $\mathrm{TopK}$, which selects only the $K$ experts with the highest scores $s_{j}({\bm{x}})$ and assigns a weight of $\infty$- ∞ to the remaining experts. Formally, the $\mathrm{TopK}$ function is defined as:
$\mathrm{TopK}\left(\bm{v},K\right)_{i}=\left\{\begin{array}[]{cc}\bm{v}_{i},&\text{ if }\bm{v}_{i}\text{ is in the }K\text{ largest elements of}\bm{v}\\ -\infty,&\text{ otherwise. }\end{array}\right.$
After the gating function selects the top $K$ experts, these experts independently compute their outputs, which are then linearly combined via their normalized scores:
@@
where $s({\bm{x}})=(s_{1}({\bm{x}}),\dots,s_{N}({\bm{x}}))$. MoE has attracted considerable attention due to their flexibility and adaptability across various domains, including large language models (Du et al., 2022; Zhou et al., 2023; Li et al., 2024), computer vision (Riquelme et al., 2021), and multi-task learning (Ma et al., 2018).
@@ 3 From Mixture of Experts to Prompt-based Continual Learning
In this section, we discuss the relationship between mixture of experts and prompt-based techniques, and demonstrate how existing prompt-based continual learning methods can be interpreted within this framework. Specifically, recent work (Le et al., 2024a, 2025c) has shown that each output vector from each attention head in an MSA layer can be viewed as the output of an MoE model. This perspective implies that an MSA layer can be regarded as a special architecture in which each attention head comprises multiple MoE models. Furthermore, these studies suggest that prefix-tuning can be interpreted as introducing new experts to fine-tune these MoE models.
Specifically, let ${\bm{X}}=\left[{\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{N}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ denote the concatenation of all input token embeddings. We define $E_{j}\in\mathbb{R}^{d\times Nd}$ as a standard basis selector such that $E_{j}{\bm{X}}={\bm{x}}_{j}$, thereby extracting the $j$-th token from the input sequence. From Equation (2), the output of the $l$-th attention head is ${\bm{h}}_{l}=\left[{\bm{h}}_{l,1},\dots,{\bm{h}}_{l,N}\right]^{\top}\in\mathbb{R}^{N\times d_{v}}$. Each output vector ${\bm{h}}_{l,i}\in\mathbb{R}^{d_{v}}$ can then be expressed as an MoE model:
${\bm{h}}_{l,i}$ $=\sum^{N}_{j=1}\frac{\exp{s_{i,j}({\bm{X}})}}{\sum^{N}_{k=1}\exp{s_{i,k}}({\bm{X}})}\cdot f_{j}({\bm{X}}),$ (5)
where the expert $f_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$, and the score function $s_{i,j}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are given by:
$f_{j}(\bm{X})$ $=W^{V^{\top}}_{l}E_{j}\bm{X}=W^{V^{\top}}_{l}\bm{x}_{j},$ (6)
for $i,j=1,\dots,N$. This structure is analogous to the Multi-gate Mixture of Experts framework (Ma et al., 2018), where multiple MoE models share a common set of experts but use distinct gating functions, as illustrated in Figure 3.
Figure 3: Illustration of the connection between attention and mixture of experts. Each attention head can be viewed as consisting of multiple MoE modules that share a common set of experts but employ different gating functions. This design closely parallels the multi-gate MoE architecture.
When the model is pre-trained, all weights $W^{Q}_{l}$, $W^{K}_{l}$, and $W^{V}_{l}$ were previously learned and then, they are held fixed, so the parameters of the experts $f_{j}$ and score functions $s_{i,j}$ remain frozen. Consequently, these pre-trained experts $f_{1},\dots,f_{N}$ can be regarded as a knowledge base. Prefix-tuning augments this knowledge base by introducing additional prefix experts encoded in prompt tokens. Specifically, the new experts and score functions are defined as follows:
$f_{N+j^{\prime}}(\bm{X})$ $=W_{l}^{V^{\top}}\bm{p}_{j^{\prime}},$ (8)
$s_{i,N+j^{\prime}}({\bm{X}})$ $=\frac{\bm{X}^{\top}E_{i}^{\top}W_{l}^{Q}W_{l}^{K^{\top}}{\bm{p}}_{j^{\prime}}}{\sqrt{d_{v}}}=\frac{{\bm{x}}_{i}^{\top}W_{l}^{Q}W_{l}^{K^{\top}}{\bm{p}}_{j^{\prime}}}{\sqrt{d_{v}}},$ (9)
for $j^{\prime}=1,\dots,L$ and $i=1,\dots,N$. From Equation (4), the output of the $l$-th attention head modified by prefix-tuning can be written as:
@@
where each output vector ${\tilde{{\bm{h}}}}_{l,i}\in\mathbb{R}^{d_{v}}$ can be demonstrated as an MoE model:
@@
for $i=1,\dots,N$. Hence, prefix-tuning can be interpreted as adding new experts $f_{N+1},\dots,f_{N+L}$ to the original pre-trained experts $f_{1},\dots,f_{N}$. These additional experts collaborate with the original ones to adapt the model for downstream tasks, alleviating the need to retrain all parameters of the model. This perspective offers a new way to analyze and explore existing prompt-based continual learning methods. Here we take L2P (Wang et al., 2022b) as a concrete example.
L2P introduces the concept of a Prompt Pool, consisting of multiple prompts that guide the pre-trained model to adapt to all tasks. Specifically, let the pool of $M$ prompts be
@@
where $\bm{K}_{\bm{x}}$ is the subset of indices for the top-$K$ keys most relevant to $\bm{x}$. The corresponding prompts are then prepended to ${\bm{x}}$ to form the prompted input ${\bm{x}}_{p}$, which is finally fed back into the Transformer model to generate the final prediction. By serving as a repository of task-specific knowledge, the prompt pool removes the need for a rehearsal buffer to prevent catastrophic forgetting (see Figure 2, right).
@@
The prompt pool can be viewed as a collection of $M\cdot L$ experts $f_{N+1},\dots,f_{N+ML}$. Each prompt ${\bm{P}}_{i}$ implicitly encodes a group of $L$ experts, $f_{N+(i-1)L+1},\dots,f_{N+iL}$ for $i=1,\dots,M$. Consequently, selecting $K$ prompts from the pool is analogous to selecting $K\times L$ experts. Each expert is assigned based on the query feature $q({\bm{x}})$, which captures contextual information. However, because each prompt ${\bm{P}}_{i}$ is associated with a single prompt key ${\bm{k}}_{i}$, all $L$ experts within a given prompt is tied to a single prompt key ${\bm{k}}_{i}$, potentially limiting the model’s expressiveness. Moreover, employing a single prompt pool for all tasks can lead to issues like forgetting or noisy knowledge, as instances from different tasks may share certain experts, thereby reducing cross-task variability.
@@
DualPrompt (Wang et al., 2022a) enhances L2P by using two complementary prompts: a general prompt (G-Prompt) and a task-specific expert prompt (E-Prompt) per task. When a new task arrives, only the G-Prompt and the newly introduced E-Prompt for that task are updated during training, while previously encountered E-Prompts remain fixed. The set of E-Prompts serves as an expanding pool of task-specific knowledge, similar to the L2P prompt pool, but with the key difference of growing incrementally with each new task. DualPrompt retains L2P’s prompt selection mechanism for E-Prompts. In contrast, G-Prompt is shared between all tasks, which requires no prompt selection. We posit that the G-Prompt aims to augment the set of pre-trained experts $f_{1},\dots,f_{N}$ by introducing additional experts capturing generalizable knowledge. However, in contrast to pre-trained experts, these newly learned experts in the G-Prompt are continually updated across tasks, thereby making them susceptible to forgetting.
@@
HiDe-Prompt (Wang et al., 2024), by comparison, uses only task-specific E-Prompts. For prompt selection, it leverages an auxiliary MLP head trained with cross-entropy loss to select the appropriate prompt, treating each task as a separate class. This class-based approach may be suboptimal because classes are defined by task order rather than semantic distinctions. Moreover, restricting each task to a single prompt may be insufficient to capture its full complexity, given the limited expressiveness of prefix-tuning (Petrov et al., 2023). Specifically, prefix experts represent offset vectors encoded by prompt tokens, rather than the linear functions used by pre-trained experts, as indicated in Equation (6) and Equation (8). This simplicity limits their capacity to accommodate the full range of variations within each task. We address these concerns in Section 4.
@@ 4 Proposed Method
In this section, we present our proposed method WAVE++ in detail. First, in Section 4.1, we examine the relationship between the mixture of experts and prefix-tuning to develop task-specific prompt pools that capture the inherent variations within individual tasks. Next, in Section 4.2, we introduce label descriptions combined with a contrastive loss to enhance the robustness of each prompt pool by encapsulating the essential characteristics of task relations. These two components collectively improve within-task performance and contribute to enhanced continual learning, as demonstrated in Equation (1). Lastly, we employ cascade voting for task prediction (Section 4.3) and generative models to mitigate classification bias toward previously learned tasks in the relation classifier (Section 4.4). An overview of our complete framework is provided in Figure 1.
@@ 4.1 Task-specific Prompt Pool
As discussed in Section 3, recent advances in continual learning, such as HiDe-Prompt (Wang et al., 2024), typically employ a dedicated prompt for each task. Although these task-specific prompts help mitigate catastrophic forgetting by leveraging specialized features, relying on a single prompt for all instances within a task can be overly restrictive. In particular, these prefix experts are simply constant functions, rendering them invariant to input variation. We posit that this limited expressiveness hinders the model from capturing the full complexity and variability inherent in each task.
@@
A straightforward strategy to enhance the expressiveness of prefix experts is to replace the constant function with a more complex one, such as a linear function. However, a key advantage of the original design lies in its simplicity and cost-effectiveness. A naive implementation of a more expressive prefix would therefore risk significantly increasing computational overhead, thus negating the practical benefits of the original approach. An alternative approach, exemplified by L2P (Wang et al., 2022b), involves maintaining a prompt pool from which prompts are selected based on the query features $q({\bm{x}})$. In this way, the set of experts is adaptively chosen according to the input, yielding a configuration similar to an SMoE architecture. This design achieves greater expressiveness while retaining efficiency.
Despite these advantages, repeatedly using the same prompts across different tasks increases the likelihood of overlap between them. Such overlap introduces the risk of overwriting task-specific knowledge, potentially erasing important information tailored to each task. Consequently, prompts may fail to capture the essential knowledge or requirements of individual tasks, leading to suboptimal within-task performance. To address this challenge, we propose constructing a prompt pool for each task, ensuring that each task maintains its distinct prompt allocation, thus reducing the risk of undesired overlap.
Specifically, for the $t$-th task, following the formulation presented in Equation (12), we introduce a prompt pool $\mathcal{P}_{t}$ consisting of $M$ prompts, defined as follows:
$\mathcal{P}_{t}$ =$=\left\{({\bm{k}}^{(t)}_{1},{\bm{P}}^{(t)}_{1}),\dots,({\bm{k}}^{(t)}_{M},{\bm{P}}^{(t)}_{M})\right\}.$ (14)
@@
Each prompt ${\bm{P}}^{(t)}_{i}\in\mathbb{R}^{L\times d}$ is associated with a learnable key ${\bm{k}}^{(t)}_{i}\in\mathbb{R}^{d}$. For prompt selection, we employ the same query-key mechanism as in L2P. We use cosine similarity for the scoring function $\gamma:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ to measure the distance between the query feature $q({\bm{x}})$, encoded using a pre-trained BERT model, and the prompt keys. The top-$K$ most relevant prompts in the prompt pool $\mathcal{P}_{t}$ are selected by optimizing the following objective:
$\bm{K}_{{\bm{x}}}$ $=\underset{S\subseteq\{1,...,M\}:|S|=K}{\rm argmin}\sum_{s\in S}\gamma(q({\bm{x}}),{\bm{k}}^{(t)}_{s}),$ (15)
where $\bm{K}_{x}$ is the set of prompt selected for the instance ${\bm{x}}$. Additionally, we introduce a prompt pool loss, as in L2P, to facilitate prompt selection:
$\mathcal{L}_{pp}=\sum_{s\in\bm{K}_{\bm{x}}}\gamma(q({\bm{x}}),{\bm{k}}_{s}^{(t)}).$ (16)
However, incorporating a prompt pool for each task may significantly increase computational costs, as each task uses multiple prompts. To address this, we configure each prompt with a single prefix expert by setting $L=1$, thereby reducing the memory cost. This approach mirrors other methods that employ a single prompt per task but with greater length. Additionally, using longer prompts would introduce multiple experts per prompt, each sharing a common prompt key ${\bm{k}}^{(t)}_{i}$. Our configuration reduces memory overhead while enhancing flexibility in expert selection during both training and testing. It allows each expert to specialize based on the input, offering a more adaptable assignment compared to methods that rely on multiple experts per prompt.
@@
Our architecture enables the assignment of distinct sets of experts to specific regions of the input data, guided by the contextual query feature $q({\bm{x}})$. This design allows each prompt in the pool to focus on the relevant information and patterns crucial for optimal performance in different regions of the input domain. As a result, the model effectively captures within-task variations. Moreover, by utilizing a task-specific prompt pool, we minimize the need for parameter sharing, thereby maximizing cross-task variance and improving overall task performance.
Comparison to Sparse Mixture of Experts
We demonstrate that our proposed task-specific prompt pool, or more generally, the prompt pool, shares several similarities with the SMoE architecture discussed in Section 2.3. Specifically, from Equation (14), by setting $L=1$, where each prompt encodes only a single prefix expert, the experts encoded in the prompt pool $\mathcal{P}_{t}$ are denoted as $f_{N+1}^{(t)},\dots,f_{N+M}^{(t)}$. Unlike the pre-trained experts $f_{1},\dots,f_{N}$ in the pre-trained BERT, which are selected by default, we employ sparse selection exclusively for these newly introduced prefix experts. As illustrated in Equation (11), each head in the MSA layer, when applying prefix-tuning, encompasses $N$ MoE models ${\tilde{{\bm{h}}}}_{l,1},\dots,{\tilde{{\bm{h}}}}_{l,N}$. A naive approach would involve applying the $\mathrm{TopK}$ function to each of these $N$ models individually, requiring the computation of all $N\times M$ score functions $s_{i,N+j^{\prime}}({\bm{X}})$ for $i=1,\dots,N$ and $j^{\prime}=1,\dots,M$. This can lead to a distinct set of prefix experts selected for each model. In contrast, our strategy utilizes the same set of $K$ new experts across all $N$ MoE models using auxiliary score functions defined as:
where $i=1,\dots,N$ and $j^{\prime}=1,\dots,M$. This strategy requires the computation of only $M$ score functions, as $\tilde{s}_{i,N+j^{\prime}}({\bm{X}})$ depends solely on $\bm{k}_{j^{\prime}}^{(t)}$, thereby enabling efficient and effective selection of $K$ experts from the prompt pool. Although the calculation of $q({\bm{x}})$might seem like an additional expense, this value is already computed for the task predictor in Section 4.3 during both the training and testing phases. As a result, the computation of $q({\bm{x}})$ can be reused without incurring any extra cost in the prompt selection process.
4.2 Enhancing Training with Label Descriptions
Let the embeddings of the input sequence ${\bm{x}}$ be denoted as ${\bm{x}}_{e}$, and the input embedding after incorporating prompts via prefix-tuning as ${\bm{x}}_{p}$. The model’s output is then given by $g_{\phi}(f_{\theta}({\bm{x}}_{p}))$, where $f_{\theta}$ represents the pre-trained transformer encoder and $g_{\phi}$ is the final relation classifier, parameterized by $\phi$.
In our framework, each task is associated with a distinct prompt pool. At test time, determining which prompt pool to apply for a given sample can be prone to errors, potentially leading to the use of an inappropriate prompt pool. Such misalignment may result in suboptimal or incorrect predictions. To mitigate this risk, we seek to improve the global generalization capacity of each prompt pool, anticipating that it may encounter samples from tasks it was not originally designed for.
To achieve this, we propose incorporating label-description representations into the training objective for each prompt pool as an auxiliary resource. These representations capture the core attributes of each relation, providing global information that helps reduce overfitting and improve model robustness. Previous studies have demonstrated that label descriptions can significantly benefit CRE tasks by providing additional context on relation types (Yang et al., 2020; Liu et al., 2022; Borchert et al., 2024; Luo et al., 2024; Thanh et al., 2025). Notably, they offer a consistent, class-specific global perspective, enabling the model to focus on essential relational characteristics and avoid spurious features.
Specifically, we leverage the detailed definitions provided for each label, typically available in benchmarking datasets, which we refer to as Raw descriptions. However, relying solely on these raw descriptions may not offer sufficient diversity and richness, leading to potential noise and instability, which could harm model performance. To address these limitations, we follow prior work (Thanh et al., 2025) and employ Gemini 1.5 (Team et al., 2023, 2024) to generate $D$ diverse, detailed, and illustrative descriptions for each relation. For each label, the raw description is fed into the large language model (LLM), serving as an expert-in-the-loop to guide the model. A depiction of our prompt template is shown in Table 1.

### v2

@@ 2.3 Mixture of Experts
Mixture of Experts (MoE) is a classical ensemble learning framework that combines multiple models to achieve more expressive and accurate predictions (Jacobs et al., 1991; Jordan and Jacobs, 1994). Formally, an MoE model consists of $N$ expert networks, denoted by $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$ for $i=1,\dots,N$. A gating function, $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$, dynamically determines the contribution of each expert for a given input ${\bm{x}}\in\mathbb{R}^{d}$. The gating function is defined by a set of learned score functions, $s_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, associated with each expert. The final output is formulated as:
${\mathbf{y}}$ $=\sum_{j=1}^{N}G({\bm{x}})_{j}\cdot f_{j}({\bm{x}})=\sum_{j=1}^{N}\frac{\exp\left(s_{j}({\bm{x}})\right)}{\sum_{\ell=1}^{N}\exp\left(s_{\ell}({\bm{x}})\right)}\cdot f_{j}({\bm{x}}).$
Building on this concept, Shazeer et al. (2017) proposed the Sparse Mixture of Experts (SMoE) architecture to efficiently scale large models. This is achieved by using a sparse gating function $\mathrm{TopK}$, which selects the $K$ experts with the highest scores $s_{j}({\bm{x}})$ and assigns a score of -$\infty$ to the remaining NKN-K experts. Formally, the $\mathrm{TopK}$ function is defined as:
[…]
After the gating function selects the top $K$ experts, their outputs are linearly combined using their normalized scores:
@@
where $s({\bm{x}})=(s_{1}({\bm{x}}),\dots,s_{N}({\bm{x}}))$. MoE models have attracted considerable attention due to their flexibility and adaptability across various domains, including large language models (Du et al., 2022; Zhou et al., 2023; Li et al., 2024), computer vision (Riquelme et al., 2021), and multi-task learning (Ma et al., 2018).
Figure 3: Illustration of the connection between attention and mixture of experts. An attention head can be interpreted as a collection of MoE modules that share a common set of experts but employ distinct gating functions. This structure closely parallels the multi-gate MoE architecture.
@@ 3 From Mixture of Experts to Prompt-based Continual Learning
In this section, we discuss the relationship between mixture of experts and prompt-based techniques, demonstrating how existing prompt-based continual learning methods can be interpreted within this framework. Specifically, recent work (Le et al., 2024a) has shown that each output vector from each attention head in an MSA layer can be viewed as the output of an MoE model. This perspective implies that an MSA layer can be regarded as a special architecture in which each attention head comprises multiple MoE models. Furthermore, these studies suggest that prefix-tuning can be interpreted as a method of fine-tuning these models by introducing new experts.
Proposition 3.1.
Let ${\bm{X}}=\left[{\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{N}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ denote the concatenation of all input token embeddings. From Equation (2), the output of the $l$-th attention head is ${\bm{h}}_{l}=\left[{\bm{h}}_{l,1},\dots,{\bm{h}}_{l,N}\right]^{\top}\in\mathbb{R}^{N\times d_{v}}$. Each output vector ${\bm{h}}_{l,i}\in\mathbb{R}^{d_{v}}$ can then be expressed as an MoE model:
${\bm{h}}_{l,i}$ $=\sum^{N}_{j=1}\frac{\exp(s_{i,j}({\bm{X}}))}{\sum^{N}_{k=1}\exp({s_{i,k}}({\bm{X}}))}\cdot f_{j}({\bm{X}}),$ (4)
where the expert function is $f_{j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$, and the score function is $s_{i,j}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}$. Both functions take the full input sequence $\bm{X}$} as input. For more details, please refer to A.1.
This structure is analogous to the Multi-gate Mixture of Experts framework (Ma et al., 2018), where multiple MoE models share a common set of experts but use distinct gating functions, as illustrated in Figure 3. In a pre-trained model, all the parameters of the experts $f_{j}$ and score functions $s_{i,j}$ are frozen. Consequently, these pre-trained experts $f_{1},\dots,f_{N}$ can be considered a knowledge base. Prefix-tuning augments this knowledge base by introducing additional prefix experts encoded as prompt tokens.
Proposition 3.2.
From Equation (3), the output of the $l$-th attention head modified by prefix-tuning can be written as:
@@
where each output vector ${\tilde{{\bm{h}}}}_{l,i}\in\mathbb{R}^{d_{v}}$ can be expressed as an MoE model:
@@
for $i=1,\dots,N$. Here, the new expert functions $f_{N+j^{\prime}}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$, and score functions $s_{i,N+j^{\prime}}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}$ are functions of $\bm{X}$} with parameters derived from ${\mathbf{P}}^{K}$ and ${\mathbf{P}}^{V}$. For more details, please refer to A.2.
Hence, prefix-tuning can be interpreted as adding new experts $f_{N+1},\dots,f_{N+L}$ to the original pre-trained experts $f_{1},\dots,f_{N}$. These additional experts collaborate with the original ones to adapt the model for downstream tasks, thus alleviating the need to retrain all of its parameters. This perspective offers a new avenue for analyzing and exploring existing prompt-based continual learning methods.
Here, we take L2P (Wang et al., 2022b) as a concrete example. L2P introduces the concept of a Prompt Pool, consisting of multiple prompts that guide the pre-trained model to adapt to all tasks. Specifically, let the pool of $M$ prompts be
@@
where $\bm{K}_{\bm{x}}$ is the set of indices for the top-$K$ keys most relevant to $\bm{x}$. The corresponding prompts are then prepended to $\bm{x}$} to form the prompted input ${\bm{x}}_{p}$, which is finally fed into the transformer model to generate the final prediction. By serving as a repository of task-specific knowledge, the prompt pool removes the need for a rehearsal buffer to prevent catastrophic forgetting (see Figure 2, right).
@@
The prompt pool can be viewed as a collection of $M\cdot L$ experts $f_{N+1},\dots,f_{N+ML}$. Each prompt ${\bm{P}}_{i}$ implicitly encodes a group of $L$ experts, $f_{N+(i-1)L+1},\dots,f_{N+iL}$ for $i=1,\dots,M$. Consequently, selecting $K$ prompts from the pool is analogous to selecting $K\times L$ experts. Each expert is assigned based on the query feature $q({\bm{x}})$, which captures contextual information. However, because each prompt ${\bm{P}}_{i}$ is associated with a single prompt key ${\bm{k}}_{i}$, all $L$ experts within a given prompt are tied to a single prompt key ${\bm{k}}_{i}$, potentially limiting the model’s expressiveness. Moreover, employing a single prompt pool for all tasks can lead to issues such as forgetting or noisy knowledge, as instances from different tasks may share certain experts, thereby reducing cross-task variability.
@@
DualPrompt (Wang et al., 2022a) enhances L2P by employing two complementary prompts: a general prompt (G-Prompt) and a task-specific expert prompt (E-Prompt) for each task. When a new task arrives, only the G-Prompt and the newly introduced E-Prompt for that task are updated during training, while E-Prompts from previous tasks remain fixed. The set of E-Prompts serves as an expanding pool of task-specific knowledge, similar to the L2P prompt pool, with the key difference that it grows incrementally with each new task. DualPrompt retains L2P’s prompt selection mechanism for E-Prompts. In contrast, the G-Prompt is shared among all tasks and thus requires no selection. We posit that the G-Prompt aims to augment the set of pre-trained experts $f_{1},\dots,f_{N}$ by introducing additional experts that capture generalizable knowledge. However, unlike the pre-trained experts, these newly learned experts in the G-Prompt are continually updated across tasks, which makes them susceptible to forgetting.
@@
HiDe-Prompt (Wang et al., 2024), by comparison, uses only task-specific E-Prompts. For prompt selection, it leverages an auxiliary MLP head trained with a cross-entropy loss to select the appropriate prompt, treating each task as a separate class. This class-based approach may be suboptimal because the classes are defined by task order rather than semantic distinctions. Moreover, restricting each task to a single prompt may be insufficient to capture its full complexity, given the limited expressiveness of prefix-tuning (Petrov et al., 2023). Specifically, prefix experts represent offset vectors encoded by prompt tokens, rather than the linear functions used by pre-trained experts, as indicated in Equation (20) and Equation (23). This simplicity limits their capacity to accommodate the full range of variations within each task. We address these concerns in Section 4.
@@ 4 Proposed Method
In this section, we present our proposed method WAVE++ in detail. First, in Section 4.1, we examine the relationship between the mixture of experts and prefix-tuning to develop task-specific prompt pools that capture the inherent variations within individual tasks. Next, in Section 4.2, we introduce label descriptions combined with a contrastive loss to enhance the robustness of each prompt pool by encapsulating the essential characteristics of task relations. These two components collectively improve within-task performance and contribute to enhanced continual learning, as demonstrated in Equation (1). Lastly, we employ cascade voting for task prediction (Section 4.3) and use generative models to mitigate classification bias toward previously learned tasks in the relation classifier (Section 4.4). Figure 1 provides an overview of our complete framework.
@@ 4.1 Task-specific Prompt Pool
As discussed in Section 3, recent advances in continual learning, such as HiDe-Prompt (Wang et al., 2024), typically employ a dedicated prompt for each task. Although these task-specific prompts help mitigate catastrophic forgetting by leveraging specialized features, relying on a single prompt for all instances within a task can be overly restrictive. In particular, these prefix experts are simply constant functions, rendering them invariant to input variations. We posit that this limited expressiveness hinders the model from capturing the full complexity and variability inherent in each task.
@@
A straightforward strategy to enhance the expressiveness of prefix experts is to replace the constant function with a more complex one, such as a linear function. However, a key advantage of the original design is its simplicity and cost-effectiveness. Naively implementing a more expressive prefix could significantly increase computational overhead, thereby undermining this core benefit. An alternative approach, exemplified by L2P (Wang et al., 2022b), involves maintaining a prompt pool from which prompts are selected based on the query features $q({\bm{x}})$. This method adaptively chooses a set of experts based on the input, yielding a configuration that resembles a SMoE architecture. This design achieves greater expressiveness while retaining computational efficiency.
Despite these advantages, repeatedly using the same prompts across different tasks increases the risk of interference. Such overlap can overwrite task-specific knowledge, erasing important information. To address this, we propose constructing a dedicated prompt pool for each task. This design isolates prompts on a per-task basis, thereby mitigating undesired overlap. Specifically, for the $t$-th task, following the formulation presented in Equation (7), we introduce a prompt pool $\mathcal{P}_{t}$ consisting of $M$ prompts, defined as follows:
$\mathcal{P}_{t}$ $=\left\{({\bm{k}}^{(t)}_{1},{\bm{P}}^{(t)}_{1}),\dots,({\bm{k}}^{(t)}_{M},{\bm{P}}^{(t)}_{M})\right\}$. (9)
@@
For prompt selection, we employ the same query-key mechanism as in L2P. We use cosine similarity for the scoring function $\gamma:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ to measure the similarity between the query feature $q({\bm{x}})$, encoded by a pre-trained BERT model, and the prompt keys. Additionally, we introduce a prompt pool loss, as in L2P, to facilitate prompt selection:
$\mathcal{L}_{pp}=\sum_{s\in\bm{K}_{\bm{x}}}\gamma(q({\bm{x}}),{\bm{k}}_{s}^{(t)}),$ (10)
where $\bm{K}_{x}$ is the set of prompts selected for the instance $\bm{x}$}.
However, maintaining a separate prompt pool for each task significantly increases computational costs. To address this issue, we configure each prompt to be a single prefix expert by setting its length to $L=1$. This approach contrasts with methods that use longer prompts, in which multiple experts must share a single, common key, ${\bm{k}}^{(t)}_{i}$. As explained in Section 3, each prompt ${\bm{P}}_{i}$ in the pool implicitly encodes a group of $L$ experts, $f_{N+(i-1)L+1},\dots,f_{N+iL}$ for $i=1,\dots,M$. Since each prompt ${\bm{P}}_{i}$ is tied to a single prompt key ${\bm{k}}_{i}$, all $L$ experts it encodes must share that same key. By setting $L=1$, we ensure that each expert has its own dedicated key, thereby avoiding the key-sharing mechanism inherent in prior methods such as L2P. This enhances flexibility, as each prefix expert can be selected independently based on the input, leading to a more specialized and adaptive model during both training and inference.
@@
Our architecture uses the contextual query feature $q({\bm{x}})$ to assign distinct sets of experts to different input instances. This design allows each expert in the prompt pool to specialize in the specific patterns and data characteristics most relevant to different regions of the input domain. As a result, the model effectively captures within-task variations. Moreover, by utilizing a task-specific prompt pool, we prevent parameter interference across tasks. This isolation preserves the integrity of task-specific knowledge, which is crucial for reducing catastrophic forgetting and enhancing overall performance. This mechanism of input-dependent expert selection closely mirrors a SMoE architecture discussed in Section 2.3. For more details, please refer to B.